OpenStudy (anonymous):

Could anyone please help me finding the general solution of an ODE? Thanks!! (xy'-1) lnx = 2y

OpenStudy (anonymous):

rearranging the equation we get, xy' -2y/lnx =1

OpenStudy (anonymous):

integrating facor= 1/ (lnx)^2

OpenStudy (anonymous):

solution will be y *I.F. = INEGRATION SIGN ( I.F.) DX +c

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

1 sec please

OpenStudy (anonymous):

Yeah thanks a lot ankur 504! I just got it following your rearrangement!!!! Thanks

OpenStudy (anonymous):

best of luck with ur maths

OpenStudy (anonymous):

Could someone tell me the steps of solving this system of differential equations? Thanks!!! x'= x-y+z y'= x+y-z z'= 2x-y

OpenStudy (anonymous):

is it a partial differential problem?

OpenStudy (anonymous):

i am not sure but the hint says that one of the eigenvalues of the correcsponding matrix is r1=1.

OpenStudy (anonymous):

right hand side makes a 3 X 3 matrix.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1 -1 1 1 1 -1 2 -1 0

OpenStudy (anonymous):

exactly!!!

OpenStudy (anonymous):

what are the steps then? I can figure the rest by myself if you don't mind giving me some directions.

OpenStudy (anonymous):

ok wait a min..

OpenStudy (anonymous):

we calculate the Jacobian?

OpenStudy (anonymous):

I gave up...I'm just a first year uni student. Sry...U better post your question on the left side...I think POLPAK can solve it

OpenStudy (anonymous):

wow that's pretty impressive. Thanks a lot though!!